OpenStudy (anonymous):

(BEST ANSWER GETS MEDAL!) if f(x)=3x^2 + 2 and g(x)=2x -1, evaluate (f-g)(-3)

OpenStudy (anonymous):

(f-g)(x) is defined by (f-g)(x)=f(x)-g(x)

OpenStudy (anonymous):

@shoplovepc so what do you think is the answer?

OpenStudy (anonymous):

Hiii @shoplovepc

OpenStudy (anonymous):

hmm I got -18?

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

@Princer_Jones

OpenStudy (anonymous):

no incorrect

OpenStudy (anonymous):

(f-g)(x)=f(x)-g(x)=(3x^2+2)-(2x-1)=3x^2+2-2x+1=3x^2-2x+3 Put x=-3 and get the solution @shoplovepc

OpenStudy (anonymous):

Thank you so much!!!

OpenStudy (anonymous):

36 is the answer